Determine whether the pair of lines is parallel,perpendicular,or neither. -3/2x+y=2, -6x+4y=10
Answer by sachi(548)

3/2x+y=2
or y=-3/2x+2=m1x+c so the slope=m1=-3/2
-6x+4y=10
or 4y=6x+10
or y=6/4x+10/4=3/2x+5/2 =m2x+c so the slope=m2=3/2
so m1*m2 not equal to -1 so not perpendicular
nor m1=m2 so not parallel